# PRISM: A Robust Framework for Skill-based Meta-Reinforcement Learning with Noisy Demonstrations

### v2

@@ C.1 Performance Comparison
At higher noise levels such as Noise($\sigma=0.2$), Noise($\sigma=0.3$) for Kitchen and Office, and Noise($\sigma=1.0$), Noise($\sigma=1.5$) for Maze2D and AntMaze, significant performance improvements highlight the effectiveness of skill refinement in addressing noisy demonstrations. Even with high-quality offline datasets like Expert and Noise($\sigma=0.1$) for Kitchen and Office, and Noise($\sigma=0.5$) for Maze2D and AntMaze, PRISM further improved performance by learning task-relevant skills. These learning curves align with the trends observed in the main experiments, confirming that skill refinement enhances performance across various dataset qualities and effectively adapts to the task distribution.
@@
Figure C.1: Performance comparison: Learning curves across considered environments and noise levels
C.2 Evolution of The Control Factor $\beta$
@@
Fig. C.2 illustrates the evolution of the control factor $\beta$ during the meta-train phase across all environments and noise levels. Initially, low $\beta$ values reflect reliance on offline datasets for skill learning, particularly in high-return environments like Kitchen and Office, where training starts with $\beta$ values close to zero. This approach prevents performance degradation by avoiding early dependence on lower-quality online samples, while ensuring gradual and stable changes in the skill distribution. As training progresses, the quality of online samples improves, leading to a gradual increase in $\beta$, which facilitates greater utilization of online data for skill refinement. For offline datasets with higher noise levels, $\beta$ converges to higher values. Consequently, PRISM learns increasingly effective skills as training proceeds, achieving superior performance on unseen tasks, underscoring the importance of PRISM’s ability to dynamically balance the use of offline and online data based on dataset quality.
@@
Figure C.2: The evolution of the control factor $\beta$ during the meta-train phase
Appendix D Additional Visualization Results for PRISM
D.1 Improvement in Task Representation through Skill Refinement
Fig. D.1 illustrates the effect of skill refinement on task representation through t-SNE visualizations of task embeddings $e^{\mathcal{T}}\sim q_{e}$ in the Kitchen environment ($\sigma=0.3$), with different tasks represented by distinct colors. In Fig. D.1 (a), the task encoder is trained using fixed skills directly derived from noisy demonstrations. The noisy skills obstruct the encoder’s ability to form clear task representations, making task differentiation challenging. This limitation highlights why, in SiMPL, relying on skills learned from noisy datasets can sometimes result in poorer performance compared to SPiRL, which focuses on task-specific skill learning.
Conversely, Fig. D.1 (b) presents the t-SNE visualization when the task encoder is trained during the meta-train phase with refined skills. The improved skills enable the encoder to form more distinct and task-specific representations, facilitating better task discrimination. This improvement allows the high-level policy to differentiate tasks more effectively and select optimal skills for each, thereby enhancing meta-RL performance. These findings demonstrate that the proposed skill refinement not only improves the low-level policy but also significantly enhances the task encoder’s ability to represent and distinguish tasks, contributing to overall performance improvements.
Figure D.1: t-SNE visualization of task representations in the Kitchen environment ($\sigma=0.3$): (a) Using only the noisy offline dataset, (b) Proposed PRISM trained with refined skills
D.2 Additional Visualizations of Refined Skills
@@
Here, we present additional visualization results for PRISM in the Kitchen and Maze2D environments. Fig. D.2 illustrates the results in the Kitchen environment ($\sigma=0.3$) after training is completed. On the left, the t-SNE visualization shows the skill representation $z\sim\pi_{h,\theta}$, while the right side highlights the distribution of skills corresponding to each subtask in the t-SNE map and how these skills solve subtasks over time in the Kitchen environment. In the t-SNE map, clusters of markers with the same shape but different colors indicate that identical subtasks share skills across different tasks. From the results, it is evident that the skills learned using the proposed PRISM framework are well-structured, with representations accurately divided according to subtasks. This enables the high-level policy to select appropriate skills for each subtask, effectively solving the tasks. Furthermore, while SiMPL trained on noisy data often succeeds in only one or two subtasks, PRISM progressively refines skills even in noisy environments, successfully solving most given subtasks.
@@
Fig. D.3 illustrates how the high-level policy utilizes refined skills obtained at different meta-train iterations (0.5K, 2K, and 4K) during the meta-test phase to solve a task in Maze2D ($\sigma=1.5$). When using skills trained solely on the offline dataset (meta-train iteration 0.5K), the agent failed to perform adequate exploration at meta-test iteration 0K. Even at meta-test iteration 0.5K, the noise within the skills hindered the agent’s ability to converge to the target task. In contrast, after refining the skills at meta-train iteration 2K, the agent successfully explored most of the maze during exploration, except for certain tasks in the upper-left region, and achieved all meta-test tasks by iteration 0.5K. Finally, using skills refined at meta-train iteration 4K, the agent not only explored almost the entire maze at meta-test iteration 0K but also completed all meta-test tasks by iteration 0.5K. Additionally, trajectories generated with refined skills showed significantly reduced deviations and shorter paths compared to those using noisy skills. Overall, the results in Fig. D.3 highlight the importance of PRISM’s skill refinement process in ensuring robust and efficient performance.
@@
Figure D.2: t-SNE visualization of skill representations (left) and refined skill trajectories for various subtasks (right) in Kitchen ($\sigma=0.3$). In the skill representation, marker colors denote tasks, while marker shapes indicate subtasks.
@@
Figure D.3: Illustration of trajectories of refined skills during the meta-test phase in Maze2D ($\sigma=1.5$), across various training and test iterations.
Appendix E Additional Ablation Studies
@@
In this section, we conduct additional ablation studies for Kitchen and Maze2D environments across all noise levels. These studies include component evaluation and PRISM’s skill refinement-related hyperparameters discussed in Section E.1, prioritization temperature $T$ in Section E.2, and the KLD coefficient $\lambda^{\text{kld}}_{\text{exp}}$ for the exploration policy in Section E.3.
E.1 Component Evaluation
@@
Fig. E.1 presents comprehensive results across all noise levels from the component evaluation in Section 5.5. While improvements are modest under conditions with high-quality offline datasets, such as Expert and Noise ($\sigma=0.1$) for Kitchen, and Expert and Noise ($\sigma=0.5$) for Maze2D, notable performance gains are still observed. The significant degradation observed in the absence of the exploration policy $\pi_{\text{exp}}$ highlights its crucial role in mitigating minor noise and discovering improved paths. For higher noise levels, such as Noise ($\sigma=0.2$), Noise ($\sigma=0.3$) for Kitchen, and Noise ($\sigma=1.0$), Noise ($\sigma=1.5$) for Maze2D, excluding the online buffer or prioritized skill refinement ($\mathcal{B}_{\text{on}}$ or $P_{\mathcal{B}_{\text{off}}}$) caused significant performance drops, emphasizing the importance of maximum return relabeling in PRISM. Additionally, relying solely on the online buffer without utilizing the offline dataset led to performance deterioration, demonstrating the offline dataset’s value in addressing meta-test tasks involving behaviors not available during meta-train. Beyond these specific findings, most trends align with the results discussed in the main text, further validating the effectiveness of PRISM’s components across different noise levels.
@@
Figure E.1: Component evaluation across all noise levels in Kitchen and Maze2D
E.2 Prioritization Temperature $T$
@@
The prioritization temperature $T$ regulates the balance between sampling from online and offline buffers. Fig. E.2 shows performance across different noise levels in Kitchen and Maze2D environments as $T$ varies. Low $T$ values lead to excessive sampling from high-return buffers, while high $T$ approximates uniform sampling, diminishing the prioritization effect. Both environments experienced degraded performance at $T=0.1$ and $T=2.0$, highlighting the importance of proper tuning. In the Kitchen environment (maximum return = 4), $T=1.0$ achieved the best performance across all noise levels, whereas in the Maze2D environment (maximum return = 1), $T=0.5$ was optimal. This difference occurs because environments with lower max returns exhibit smaller gaps between low- and high-return buffers, reducing the effect of prioritization. Based on these findings, we set the best-performing hyperparameter values as defaults for each environment.
E.3 KLD Coefficient $\lambda^{\text{kld}}_{\text{exp}}$
@@
The KLD coefficient $\lambda^{\text{kld}}_{\text{exp}}$ regulates the strength of the KLD term between the exploration policy and the action distribution induced by the prioritized online buffer $\mathcal{B}_{\text{on}}^{i}$ for each task $\mathcal{T}^{i}$. Fig. E.3 illustrates performance variations with $\lambda^{\text{kld}}_{\text{exp}}\in[0,0.001,0.002,0.005]$.
@@
In the Kitchen environment, $\lambda^{\text{kld}}_{\text{exp}}=0.005$ performed best for Expert and Noise ($\sigma=0.1$), while $\lambda^{\text{kld}}_{\text{exp}}=0.002$ and $\lambda^{\text{kld}}_{\text{exp}}=0.001$ were optimal for Noise ($\sigma=0.2$) and Noise ($\sigma=0.3$), respectively. At lower noise levels, the high-level policy benefits from quickly following high-return samples, whereas at higher noise levels, focusing on exploration to discover shorter paths becomes more advantageous. For the Maze2D environment, performance was consistent across $\lambda^{\text{kld}}_{\text{exp}}=0.001,0.002,$ and $0.005$, with only minor variations observed. However, when $\lambda^{\text{kld}}_{\text{exp}}=0$, removing the KLD term resulted in significant performance degradation across all noise levels in both Kitchen and Maze2D environments. This highlights the necessity of guidance from high-return samples for effectively solving long-horizon tasks. Based on these results, we selected the best-performing hyperparameter values as defaults for each environment.
@@
Figure E.2: Ablation study for the prioritization temperature $T$ across all noise levels in Kitchen and Maze2D
@@
Figure E.3: Ablation study for the KLD coefficient $\lambda^{\text{kld}}_{\text{exp}}$ across all noise levels in Kitchen and Maze2D
Appendix F Limitations
The PRISM algorithm introduces a robust skill-based meta-RL method through a novel skill refinement approach. By incorporating skill exploration, it discovers useful behaviors around demonstrations, and through maximum return relabeling, it enables prioritized skill refinement that is resilient to demonstration quality. However, there are some potential areas for improvement. First, simultaneously training both low- and high-level policies in PRISM can increase computational costs and memory usage. Nevertheless, as mentioned, the iteration count is calculated separately for exploration and high-level policy updates, keeping the computational load per iteration comparable. Additionally, memory issues can be mitigated by adjusting buffer sizes. Importantly, PRISM demonstrates superior performance even when baseline algorithms are trained for extended periods, indicating that the benefits of PRISM outweigh its drawbacks. Another limitation is that fine-tuning is required during the meta-test phase. However, this is an inherent aspect of skill-based learning. Future work could focus on enabling zero-shot skill adaptation to further enhance PRISM’s utility and applicability.

### v3

@@ D.1 Performance Comparison
Fig. D.1 presents the learning curves of average returns for the algorithms on test tasks, corresponding to the experiments summarized in Table 1. Rows represent evaluation environments, and columns denote noise levels. SISL consistently demonstrated superior robustness, outperforming all baselines across various environments and noise levels. At higher noise levels such as Noise($\sigma=0.2$), Noise($\sigma=0.3$) for Kitchen and Office, and Noise($\sigma=1.0$), Noise($\sigma=1.5$) for Maze2D and AntMaze, significant performance improvements highlight the effectiveness of skill refinement in addressing noisy demonstrations. Even with high-quality offline datasets like Expert and Noise($\sigma=0.1$) for Kitchen and Office, and Noise($\sigma=0.5$) for Maze2D and AntMaze, SISL further improved performance by learning task-relevant skills. These learning curves align with the trends observed in the main experiments, confirming that skill refinement enhances performance across various dataset qualities and effectively adapts to the task distribution.
Interestingly, SPiRL and SiMPL sometimes perform better with mild Gaussian noise than with expert data, especially in Maze2D and AntMaze. Our analysis suggests that mild noise increases the diversity of behaviors in the dataset without significantly reducing success rates, allowing agents to reach goals via more diverse paths. This expands state-action coverage (by about 7.3% in Maze2D and 5.2% in AntMaze), helping skill-based methods learn more flexible and reusable skills. However, as the noise level increases further, a significant portion of trajectories fail to reach the goal, leading to low quality skill learning and degraded downstream performance.
@@
Figure D.1: Learning curves across considered environments and noise levels
D.2 Limited Offline Dataset Size
To further investigate the robustness of SISL, we conducted additional experiments to assess the impact of dataset size on skill learning and downstream performance. While SISL primarily addresses the challenge of refining corrupted skills from noisy demonstrations through online interaction, the size of the offline dataset remain important factors, especially in practical scenarios. To assess the impact of dataset size, we conducted additional experiments on the Kitchen environment using the same expert dataset but reduced to 50% (0.5M), 25% (0.25M), and 10% (0.1M) of its original dataset size (1M). As shown in Table D.1, SISL consistently outperforms baselines even under limited expert data, and its performance degrades more gracefully compared to baselines. These results confirm SISL’s ability to refine skills even from limited data, which is particularly valuable in real-world scenarios where collecting high-quality demonstrations is costly or infeasible.
Table D.1: Final performance on the Kitchen environment with varying sizes of expert datasets.
Dataset Size SPiRL SiMPL SISL
[…]
D.3 Random Noise Injection
While Gaussian noise is a widely used approach for degrading demonstration quality in offline RL studies, it does not fully capture the diverse and unstructured nature of real-world noise such as sensor failures, occlusions, or actuator malfunctions. In our main experiments, we adopted multi-level Gaussian noise for two reasons: (1) it is a standard and accepted method for systematically degrading demonstration quality, and (2) it enables controlled analysis of robustness under varying degrees of skill degradation. Notably, in domains such as Kitchen and Office, sufficiently high levels of Gaussian noise render learned skills nearly unusable, effectively mimicking real-world failure scenarios in precision control tasks and providing a highly challenging regime for evaluating robustness.
To further assess the generality of our robustness claims and address concerns about the limitations of Gaussian noise, we conducted an additional experiment using random action injection in the Kitchen environment. This approach better simulates real-world anomalies such as actuator faults or sensor failures. Specifically, at each timestep, the oracle action was replaced with a randomly sampled action with a probability of 25%, 50%, or 100% (resulting in a uniformly random dataset at the highest level). This method introduces severe, unstructured corruption into the offline dataset, representing worst-case real-world failures beyond the smooth degradations induced by Gaussian noise. As shown in Table D.2, SISL consistently outperformed baseline methods, confirming its robustness even under extreme, non-gaussian corruption. These results demonstrate SISL’s ability to refine useful behaviors and maintain strong performance in the presence of severe dataset corruption, further validating the generality of our robustness claims.
Table D.2: Final performance on the Kitchen environment with the random noise injection.
Noise Type SPiRL SiMPL SISL
[…]
D.4 Diverse Sub-Optimal Dataset
Beyond injecting noise into expert demonstrations, we also considered datasets of diverse quality sub-optimal demonstration. Our decision to focus on noisy expert trajectories stems from the specific challenge we aim to address. Unlike offline-RL, which typically assumes access to reward-labeled trajectories and aims to improve performance from sub-optimal data, our setup considers reward-free offline data where noise degrades originally near-optimal demonstrations. This setting better reflects our goal of studying how exploration can help enhance skill learning when clean supervision is unavailable. To test whether robustness hold across different data qualities, we conducted additional experiments in the Kitchen domain using three dataset types: expert (return == 4), medium (return $\approx$ 2), and random (collected from a uniformly random policy). As summarized in Table D.3, SISL consistently outperforms other methods across all dataset types, further validating its robustness to data sub-optimality.
Table D.3: Final performance on the Kitchen environment with diverse quality of offline datasets.
Dataset SPiRL SiMPL SISL
[…]
Appendix E Computational Complexity
Despite SISL’s significant performance improvements on noisy demonstrations through self-improving skill refinement, it is important to consider the computational trade-offs. Simultaneously training both low-level and high-level policies in SISL can increase computation time. During meta-train, SISL requires approximately 16% more time per iteration (SiMPL: 15.1s vs SISL: 17.6s in Kitchen). However, as shown in Fig. 6, SiMPL does not improve under high-noise settings even with extended training, whereas SISL continues to improve, justifying this computational overhead. Importantly, considering that the main goal of meta-RL is adaptation to unseen tasks, SISL does not require additional computational cost during the fine-tuning phase and maintains similar computation time as SiMPL in the meta-test phase while still achieving superior performance compared to baseline algorithms.
Appendix F Further Analysis Results for the SISL Framework
In this section, we provide a more detailed analysis and visualization results of the skill refinement process in the SISL. The analysis performed includes the evolution of the mixing coefficient $\beta$ in Section F.1, visualization of the refined skills and the corresponding skill sequence visualization in Section F.2, task representation in Section F.3, and the learning stability of the reward model during the meta-train phase in Section F.4.
F.1 Evolution of The Mixing Coefficient $\beta$
@@
Fig. F.1 illustrates the evolution of the mixing coefficient $\beta$ during the meta-train phase across all environments and noise levels. Initially, low $\beta$ values reflect reliance on offline datasets for skill learning, particularly in high-return environments like Kitchen and Office, where training starts with $\beta$ values close to zero. This approach prevents performance degradation by avoiding early dependence on lower-quality online samples, while ensuring gradual and stable changes in the skill distribution. As training progresses, the quality of online samples improves, leading to a gradual increase in $\beta$, which facilitates greater utilization of online data for skill refinement. For offline datasets with higher noise levels, $\beta$ converges to higher values. Consequently, SISL learns increasingly effective skills as training proceeds, achieving superior performance on unseen tasks, underscoring the importance of SISL’s ability to dynamically balance the use of offline and online data based on dataset quality.
@@
Figure F.1: The evolution of the mixing coefficient $\beta$ during the meta-train phase
F.2 Additional Visualizations of Refined Skills
@@
Here, we present additional visualization results for SISL in the Kitchen and Maze2D environments. Fig. F.2 illustrates the results in the Kitchen environment ($\sigma=0.3$) after training is completed. On the left, the t-SNE visualization shows the skill representation $z\sim\pi_{h,\theta}$, while the right side highlights the distribution of skills corresponding to each subtask in the t-SNE map and how these skills solve subtasks over time in the Kitchen environment. In the t-SNE map, clusters of markers with the same shape but different colors indicate that identical subtasks share skills across different tasks. From the results, it is evident that the skills learned using the proposed SISL framework are well-structured, with representations accurately divided according to subtasks. This enables the high-level policy to select appropriate skills for each subtask, effectively solving the tasks. Furthermore, while SiMPL trained on noisy data often succeeds in only one or two subtasks, SISL progressively refines skills even in noisy environments, successfully solving most given subtasks.
@@
Fig. F.3 illustrates how the high-level policy utilizes refined skills obtained at different meta-train iterations (0.5K, 2K, and 4K) during the meta-test phase to solve a task in Maze2D ($\sigma=1.5$). When using skills trained solely on the offline dataset (meta-train iteration 0.5K), the agent failed to perform adequate exploration at meta-test iteration 0K. Even at meta-test iteration 0.5K, the noise within the skills hindered the agent’s ability to converge to the target task. In contrast, after refining the skills at meta-train iteration 2K, the agent successfully explored most of the maze during exploration, except for certain tasks in the upper-left region, and achieved all meta-test tasks by iteration 0.5K. Finally, using skills refined at meta-train iteration 4K, the agent not only explored almost the entire maze at meta-test iteration 0K but also completed all meta-test tasks by iteration 0.5K. Additionally, trajectories generated with refined skills showed significantly reduced deviations and shorter paths compared to those using noisy skills. Overall, the results in Fig. F.3 highlight the importance of SISL’s skill refinement process in ensuring robust and efficient performance.
@@
Figure F.2: t-SNE visualization of skill representations (left) and refined skill trajectories for various subtasks (right) in Kitchen ($\sigma=0.3$). In the skill representation, marker colors denote tasks, while marker shapes indicate subtasks.
@@
Figure F.3: Illustration of trajectories of refined skills during the meta-test phase in Maze2D ($\sigma=1.5$), across various training and test iterations.
F.3 Improvement in Task Representation through Skill Refinement
Fig. F.4 illustrates the effect of skill refinement on task representation through t-SNE visualizations of task embeddings $e^{\mathcal{T}}\sim q_{e}$ in the Kitchen environment ($\sigma=0.3$), with different tasks represented by distinct colors. In Fig. F.4 (a), the task encoder is trained using fixed skills directly derived from noisy demonstrations. The noisy skills obstruct the encoder’s ability to form clear task representations, making task differentiation challenging. This limitation highlights why, in SiMPL, relying on skills learned from noisy datasets can sometimes result in poorer performance compared to SPiRL, which focuses on task-specific skill learning.
Conversely, Fig. F.4 (b) presents the t-SNE visualization when the task encoder is trained during the meta-train phase with refined skills. The improved skills enable the encoder to form more distinct and task-specific representations, facilitating better task discrimination. This improvement allows the high-level policy to differentiate tasks more effectively and select optimal skills for each, thereby enhancing meta-RL performance. These findings demonstrate that the proposed skill refinement not only improves the low-level policy but also significantly enhances the task encoder’s ability to represent and distinguish tasks, contributing to overall performance improvements.
Figure F.4: t-SNE visualization of task representations in the Kitchen environment ($\sigma=0.3$):
(a) Using only the noisy offline dataset, (b) Proposed SISL trained with refined skills
F.4 Stability of Reward Model Learning
Unlike offline-RL, which typically relies on reward-labeled datasets, skill-based RL operates on reward-free offline data. In SISL, only the skills are learned from the noisy offline trajectories, while the reward model and the high-level policy are trained online using clean, noise-free training tasks, as in SiMPL. As shown in Eq. (4), the reward model is trained using transitions sampled from $\mathcal{B}^{i}_{\text{imp}}$ and $\mathcal{B}^{i}_{\text{on}}$, which contain only online trajectories collected by interacting with training task $i$. Since these buffers contain no offline data, the reward model is unaffected by offline noise. To further support this point, we present Table F.1, which shows that the MSE of the reward model remains consistently low across different offline noise settings in the Kitchen environment. These results indicate that the reward model maintains high accuracy and continues to support effective relabeling even when offline skill pretraining is noisy.
Table F.1: MSE of reward model under different noise levels in the Kitchen environment.
Kitchen(Expert) Kitchen($\sigma$=0.1) Kitchen($\sigma$=0.2) Kitchen($\sigma$=0.3)
[…]
Appendix G Additional Ablation Studies
@@
In this section, we conduct additional ablation studies for Kitchen and Maze2D environments across all noise levels. These studies include component evaluation and SISL’s skill refinement-related hyperparameters discussed in Section G.1, prioritization temperature $T$ in Section G.2, the KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ for the skill-improvement policy in Section G.3, and additional component evaluation on RND and re-initialization in Section G.4.
G.1 Component Evaluation
@@
Fig. G.1 presents comprehensive results across all noise levels from the component evaluation in Section 5.5. While improvements are modest under conditions with high-quality offline datasets, such as Expert and Noise ($\sigma=0.1$) for Kitchen, and Expert and Noise ($\sigma=0.5$) for Maze2D, notable performance gains are still observed. The significant degradation observed in the absence of the skill-improvement policy $\pi_{\text{imp}}$ highlights its crucial role in mitigating minor noise and discovering improved paths. For higher noise levels, such as Noise ($\sigma=0.2$), Noise ($\sigma=0.3$) for Kitchen, and Noise ($\sigma=1.0$), Noise ($\sigma=1.5$) for Maze2D, excluding the online buffer or skill prioritization via maximum return relabeling ($\mathcal{B}_{\text{on}}$ or $P_{\mathcal{B}_{\text{off}}}$) caused significant performance drops, emphasizing the importance of maximum return relabeling in SISL. Additionally, relying solely on the online buffer without utilizing the offline dataset led to performance deterioration, demonstrating the offline dataset’s value in addressing meta-test tasks involving behaviors not available during meta-train. Beyond these specific findings, most trends align with the results discussed in the main text, further validating the effectiveness of SISL’s components across different noise levels.
@@
Figure G.1: Component evaluation across all noise levels in Kitchen and Maze2D
G.2 Prioritization Temperature $T$
@@
The prioritization temperature $T$ regulates the balance between sampling from online and offline buffers. Fig. G.2 shows performance across different noise levels in Kitchen and Maze2D environments as $T$ varies. Low $T$ values lead to excessive sampling from high-return buffers, while high $T$ approximates uniform sampling, diminishing the prioritization effect. Both environments experienced degraded performance at $T=0.1$ and $T=2.0$, highlighting the importance of proper tuning. In the Kitchen environment (maximum return = 4), $T=1.0$ achieved the best performance across all noise levels, whereas in the Maze2D environment (maximum return = 1), $T=0.5$ was optimal. This difference occurs because environments with lower max returns exhibit smaller gaps between low- and high-return buffers, reducing the effect of prioritization. Based on these findings, we suggest a practical guideline: select $T$ roughly in proportion to the environment’s maximum achievable return. This approach offers a principled starting point for tuning $T$ without requiring an extensive hyperparameter sweep, thereby improving usability and reproducibility. Accordingly, we set the best-performing hyperparameter values as defaults for each environment.
G.3 KLD Coefficient $\lambda^{\textnormal{kld}}_{\textnormal{imp}}$
@@
The KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ regulates the strength of the KLD term between the skill-improvement policy and the action distribution induced by the prioritized online buffer $\mathcal{B}_{\text{on}}^{i}$ for each task $\mathcal{T}^{i}$. Fig. G.3 illustrates performance variations with $\lambda^{\text{kld}}_{\text{imp}}\in[0,0.001,0.002,0.005]$.
@@
In the Kitchen environment, $\lambda^{\text{kld}}_{\text{imp}}=0.005$ performed best for Expert and Noise ($\sigma=0.1$), while $\lambda^{\text{kld}}_{\text{imp}}=0.002$ and $\lambda^{\text{kld}}_{\text{imp}}=0.001$ were optimal for Noise ($\sigma=0.2$) and Noise ($\sigma=0.3$), respectively. At lower noise levels, the high-level policy benefits from quickly following high-return samples, whereas at higher noise levels, focusing on exploration to discover shorter paths becomes more advantageous. For the Maze2D environment, performance was consistent across $\lambda^{\text{kld}}_{\text{imp}}=0.001,0.002,$ and $0.005$, with only minor variations observed. However, when $\lambda^{\text{kld}}_{\text{imp}}=0$, removing the KLD term resulted in significant performance degradation across all noise levels in both Kitchen and Maze2D environments. This highlights the necessity of guidance from high-return samples for effectively solving long-horizon tasks. Based on these results, we selected the best-performing hyperparameter values as defaults for each environment.
@@
Figure G.2: Impact of the prioritization temperature $T$ across all noise levels in Kitchen and Maze2D
@@
Figure G.3: Impact of the KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ across all noise levels in Kitchen and Maze2D
G.4 Additional Component Evaluation
Following the ablation of SISL’s core components in Section G.1, we performed additional ablations on the remaining components that could affect performance. These evaluations are conducted in the Kitchen environment using both a Expert and a high-noise condition Noise($\sigma=0.3$), and the results are shown in Table G.1.
The SISL w/o RND is a variant that excludes RND from skill-improvement policy $\pi_{\text{imp}}$ training and relies solely on SAC. The results show performance degradation compared to full SISL, with especially large drops under high-noise conditions where exploration capability is critical. This suggests that RND encourages exploration toward rare or less frequently visited states, enabling the discovery of more diverse and useful trajectories.
The SISL w/o Re-Initialize refers to a variant that continues training the high-level policy $\pi_{h}$ without re-initializing it at each $K_{\text{iter}}$ during the meta-training process. In SISL, the skill model is periodically updated every $K_{\text{iter}}$ steps, which effectively changes the environment dynamics observed by the high-level policy. Continuing to train the same high-level policy across different skill sets introduces non-stationarity, leading to instability. To address this, we reset both the policy parameters and the buffer at each skill update so that the high-level policy can re-learn from scratch under a new, stable MDP defined by the updated skills. This technique is consistent with practices in continual and safe reinforcement learning, where re-initialization is often used to manage sudden changes in task dynamics (Kim et al., 2023; Kong et al., 2024). To empirically validate this decision, we conducted ablation experiments on high-level policy re-initialization. The results show that removing re-initialization leads to a substantial performance drop, supporting the necessity of this design choice.
Table G.1: Comparison of SISL with/without RND and re-initialization.
Dataset SISL SISL w/o RND SISL w/o Re-Initialize
[…]